Circular No. 5833

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

A0535+26 R. B. Wilson, M. H. Finger, L. Gibby, and G. J. Fishman report for

the BATSE/Compton GRO Observatory team: “Hard x-ray pulses in the energy range ∼ 20-120 keV are currently being detected from A0535+26. The present outburst was first detected on July 8. The pulsed flux increased slightly from then until July 13, and has remained approximately constant through July 17. The spectrum of the phase averaged pulsed flux on July 13 is adequately fit by a power law, having spectral index –2.3 ± 0.5, with a flux of 5.7 ± 1.1 × 10^{−5} photon cm^{−2} s^{−1} keV^{−1} at 50 keV. The pulse shape is very broad, with a sharp dip occupying ∼ 0.2 in phase. The pulse period on July 17 was 103.377 ± 0.005 s, corrected to the solar system barycenter. A reanalysis of earlier BATSE data shows that the source was also detected at weaker levels during Feb. 28-Mar. 26 (just prior to the optical activity reported from its companion HDE 245770 by Bartolini et al. on IAUC 5746) and June 20-27. Additional observations at optical and other wavelengths are encouraged.”

COMET MUELLER (1993a)

Ephemeris from orbital elements on MPC 22383:

1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
July 12 7^{h}21.^{m}85 +56^{o}27.^{′}9 3.705 2.924 34.^{o}5 11.^{o}4 12.0
22 7 32.41 +57 21.5 3.579 2.842 37.6 12.6 11.8
Aug. 1 7 43.99 +58 31.9 3.431 2.761 42.0 14.2 11.6
11 7 56.75 +60 02.5 3.262 2.682 47.4 16.2 11.4
21 8 11.05 +61 57.7 3.075 2.605 53.6 18.2 11.1
31 8 27.51 +64 23.5 2.874 2.529 60.2 20.3 10.8
Sept. 10 8 47.58 +67 26.8 2.665 2.457 67.2 22.2 10.5
20 9 14.6 +71 15.3 2.453 2.387 74.4 23.9 10.2
30 9 57.9 +75 52.4 2.247 2.320 81.5 25.3  9.9
Oct. 10 11 33.0 +80 50.4 2.057 2.257 88.1 26.2  9.6
20 15 28.4 +82 32.4 1.897 2.198 93.7 26.9  9.3
30 18 25.9 +76 25.1 1.781 2.144 97.1 27.4  9.1
Nov. 9 19 33.08 +66 52.4 1.720 2.095 97.6 27.9  8.9
19 20 08.17 +56 29.2 1.722 2.052 94.5 28.7  8.8
29 20 31.82 +46 30.2 1.783 2.015 88.5 29.3  8.8

1993 July 23 Daniel W. E. Green